OpenStudy (fearbigjjrob):

The graph of f(x) = 2^x + 1 is shown below. Explain how to find the average rate of change between x = 0 and x = 2.

OpenStudy (fearbigjjrob):

@hartnn can you help? I have no idea how to do this, there is also a graph im going to post that went with the question

hartnn (hartnn):

know the formula? :)

hartnn (hartnn):

or does this need to be done using graph only?

OpenStudy (fearbigjjrob):

um no.... and I actually want to work it out fully not just the answer... cause I know some people just give it, which is nice but I dont get how it works then....

OpenStudy (fearbigjjrob):

ummm I am not sure about the graph part....

hartnn (hartnn):

It can be done using a simple formula (similar to slope): \(\Large \dfrac{f(b)-f(a)}{b-a}\) here, f(x) = 2^x +1 a= 0 b =2

OpenStudy (fearbigjjrob):

ok give me a second to work that out

hartnn (hartnn):

so first get f(0) and f(2) either from graph or by plugging in

OpenStudy (fearbigjjrob):

um what do you mean by that? just plug them in, and actually I have no idea what to do.....

hartnn (hartnn):

f(x) = 2^x + 1 to get f(0) , just replace 'x' by '0' f(0) = 2^0 + 1 = ..?

OpenStudy (fearbigjjrob):

ohhhhh ok one minute

OpenStudy (fearbigjjrob):

2 to the 0 power would just be 2 correct? I have never done a 0 power before, and in that case f(0)=3

hartnn (hartnn):

(anything except 0)^0 = 1

OpenStudy (fearbigjjrob):

oh f(0)=1 not 3

hartnn (hartnn):

f(0)= 2^0+1=1+1=

OpenStudy (fearbigjjrob):

2?

hartnn (hartnn):

yes f(0)=2 how about f(2)=..?

OpenStudy (fearbigjjrob):

f(2)=5 I believe

hartnn (hartnn):

yesss, just plug these values in the formula i gave

OpenStudy (fearbigjjrob):

\[\frac{ 5-2 }{ 2-0 }=\frac{ 3 }{ -2 }\] or would the -2 make the whole thing negative?

OpenStudy (fearbigjjrob):

ignore that for a minute....

OpenStudy (fearbigjjrob):

so looking back at it I think I messed up with that, it should be \[\frac{ 3 }{ 3 }\] correct?

hartnn (hartnn):

5-2 = 3 2-0 = 2 so, the rate of change is 3/2 :)

OpenStudy (fearbigjjrob):

ohh ok, sorry I was good the first time, I originally thought the bottom ones were switched to f(b) and f(a) sorry had an idiot moment... but thanks for you help :D

hartnn (hartnn):

welcome ^_^ do you know how to find slope of line give 2 points?

OpenStudy (fearbigjjrob):

I believe you need to use slope-intercept form correct? or a formula like x1-x2/y1-y2..... its been a while since I did slope actually...

hartnn (hartnn):

yes, the formula is \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) Just wanted to let you know the 2 independent ways of doing this same problem/ 1. The formula way, which we did here. Only using f(x). 2. The graph way: >> Find the y values from graph for the 2 x values given. So, you'll get y = 2 for x = 0, y = 5 for x= 2 so, your 2 points are (0,2),(2,5) Find the slope of line containing these 2 points, and you'll get the same answer :)

OpenStudy (fearbigjjrob):

ahh ok, thanks, I am gonna write it down, im studing up cause Im gonna retake the Math PERT test to try getting into college math, and questions like this were common on it :)

hartnn (hartnn):

cool! and good luck :)